OpenStudy (anonymous):

2 – 3x = x – 10 2 – 3x = x – 10 @Mathematics

OpenStudy (anonymous):

2x = -10 - 2 x = -12/2 = -6

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

-10-2=-3x-x -12=-4x x=3